# Automatic Prompt Optimization for Knowledge Graph Construction: Insights from an Empirical Study

### v1

@@ 1 Introduction
Recent advances in large language models, which have significantly improved performance on core NLP tasks, have transformed the research area of KG construction from text [6]. Most of the early large language model (LLM)-based methods approach the problem in two key ways. The first is as an extraction of factual and common-sense knowledge using pre-trained language models via prompt-based prediction of masked objects from partial sentences describing complete triples [7]. The second is by fine-tuning an LLM on the autoregressive generation of text for triple extraction [8, 9]. Alternatives such as Re2G [10] and GraphRAG [11] use LLMs to build Retrieval Augmented Generation (RAG) systems [12].
@@ 2 Triple Extraction for Knowledge Graph Construction
To build a Knowledge Graph from text corpora, a system needs to convert the unstructured natural-language text into a set of elemental triples. Triple extraction involves sub-tasks such as identifying the core entities mentioned in the text (Named Entity Recognition or NER), along with identification/extraction of relations that constitute the facts mentioned in the text (Relation Extraction or RE). These sub-tasks can be pipelined, by performing entity extraction first, and then extracting relations (+triples) or in a joint manner, where entities, relations, and triples are extracted simultaneously (end-to-end relation extraction). In this paper, we focus on end-to-end triple extraction, where the input is a natural language text and the expected output is a set of triples.
@@ 3 Automated Prompt Optimization
Gradient-based approaches such as ProTeGI [32], MAPO [33], TextGrad [34], etc. use gradient descent-like algorithms to optimize prompt embeddings according to a predefined performance objective [35]. Starting from a human prompt, these methods learn continuous valued embeddings using the gradient directions. On the other hand, approaches such as InstructZero [36] optimize low-dimensional “soft prompts" using LLMs instead of directly optimizing discrete instructions. They balance exploitation vs exploration in this space using Bayesian Optimization techniques. In terms of discrete optimization, approaches such as EvoPrompt [37] circumvent access to gradients and internal parameters of LLMs using Evolutionary Algorithms to generate candidate prompts, thus making them more general. Another approach to the same problem leverages in-context learning. For example, Automated Prompt Engineer (APE) [38] generates prompts based on a few real-life examples, tailoring them to the task without needing extensive training data. APE treats the instruction as a program and optimizes it by searching through a pool of instruction candidates proposed by an LLM to maximize a chosen score function. A complementary approach to APO uses reinforcement learning-based strategies such as OIRL [39], which model the interaction between the query-prompt pair via a reward model for proposing and evaluating candidate prompts suited for arithmetic reasoning tasks. In contrast, meta-Prompting [40] uses structural and syntactical aspects of tasks to create general prompts that guide the generation of task-specific prompts. In lieu of template-based approaches, the work of [41] called DSPy frames prompt engineering as a declarative, compiler-driven optimization task. DSPy poses prompting as a computational graph pipeline for text transformation, where LLMs are invoked through parameterized self-improving modules. This provides great flexibility across diverse tasks given a computational budget.
@@ 4.1 Task
Let $\mathbf{x}=(x_{1},x_{2},\dots,x_{n})$ denote a passage of natural language text with a list of terms, and let =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{m}\}$ be a predefined set of canonical relation types. The goal of the triple extraction task is to identify: (a) A set of entity mentions =$\mathcal{E}=\{e_{1},e_{2},\dots,e_{k}\}$ from the text, (b) a subset of relations $r\in\mathcal{R}$ expressed in the text, and (c) a set of directed triples $\bigl{\{}(e_{i},\,r,\,e_{j})\mid e_{i},e_{j}\in\mathcal{E},\;r\in\mathcal{R}\bigr{\}}$ that are represents the semantic relationships between the entity pairs that are mentioned in text. An example of the task, along with inputs and outputs, is illustrated in Figure 1.
@@ 4.3 Automatic Prompt Optimizers
DSPy [41] (Declarative Self-improving Python) is an open-source toolkit for modular prompt design as programming with structured and declarative natural-language modules. DSPy allows specifying input/output behaviour as a signature and selecting a module (e.g., predict, Chain-Of-Thought, ReAct) to assign a strategy for invoking the LLM. DSPy natives parse the LLM output based on the pre-defined signature. It allows for the optimization of prompts using a training set, which is analyzed to create a ‘sub-dataset’ and task descriptions, propose better prompts with LLM-generated instructions, and select effective few-shots examples. Specifically, we used MIPROv2 (Multiprompt Instruction PRoposal Optimizer Version 2) [44] to optimize instructions and few-shot examples simultaneously as a joint search problem. It first bootstraps a set of few-shot example candidates, proposes instructions grounded in different task dynamics, and finds an optimized combination of these joint candidates using Bayesian optimization. Figure 2 shows an example of an automatically optimized prompt using DSPy.
@@
APE [38] APE does not require any initial prompt. Given some examples of input and expected output, it generates several task instruction candidates, executes them using the target model, and selects the most appropriate instruction based on evaluation scores. By default, APE uses GPT-3 (text-davinci-002 to be more specific). We replace it with Granite-3.2-8B-Instruct. We used the forward generation template settings and randomly picked 10 examples from the SynthIE validation data to provide as few shots, and selected the generated prompt with the highest score based on the validation set of SynthIE.
@@ 5.1 Prompting strategy
Table 2 outlines the performance improvements in terms of macro Precision, Recall, and F1 obtained by prompt optimization. We observe that all prompting strategies result in a performance gain, as seen in Table 2. Of these, Chain-Of-Thought asked to solely extract triples obtains the best triple extraction F1 of $0.73$. The highest overall performance improvement is seen in the case of the Extract, Critique, and Refine pipeline run on entities, relations, and triples i.e., ECR (E-R-T) output, with a $+16\%$ improvement in triple extraction F1.
@@ 5.3 Schema Complexity
To analyze this research question, we change the number of allowed relations that we include in the prompt. When a larger number of relations are included in the input text, we expect that the increased search space makes the task slightly harder for the systems with the increase in relation types. We examine eight settings with the number of allowed relation types ranging from $100-800$.
@@ 5.5 Train, Validation and Test Datasets
As shown in Table 6, the results demonstrate that using the same dataset for both few-shot selection and prompt validation significantly outperforms the case when distinct datasets are used for training/validation and testing. For instance, when SynthIE is used as the test set, the Predict (E-R-T) prompt improves triple F1 by $+0.01$ when trained on the REBEL dataset, compared to a $+0.08$ improvement when the same dataset is used for both training and testing. Similarly, with Predict (T) and SynthIE as the test set, using REBEL for training/validation negatively impacts the optimized prompt, reducing triple F1 by $-0.04$ points. When REBEL is used as the test set, a similar trend is observed, with a larger margin of improvement when the same dataset is used for both training and testing compared to when different datasets are employed. Our hypothesis is that using few-shot examples and dataset summaries generated from a dataset that differs from the final test introduces some bias, leading to a prompt that is suboptimal for the test set.
@@ 6 Qualitative Analysis of Generated Prompts
Interesting, both the prompts yieleded similar results (when fed into the Llama3.3-70B) on the validation data - F1-Entities: $0.95$/F1-Relations: $0.77$/F1-Triples: $0.70$.
@@ 7 Conclusions
In this empirical study, we explored the application of automatic prompt optimization for triple extraction via extensive experimental benchmarking. We evaluated different settings by changing (a) the prompting strategy, (b) the LLM being used for prompt optimization and task execution, (c) number of canonical relations in the schema (schema complexity), (d) the length and diversity of input text, (e) the metric used to drive the prompt optimization, and (f) the dataset being used for training and testing. Based on the empirical results, we can conclude that automatic prompt optimization can be used to generate reasonable prompts for the triple extraction task with improved performance compared to baseline prompts. All three prompt optimization approaches that were tested, DSPy, APE, and TextGrad, consistently outperformed the baseline prompt in precision, recall, and F1 metrics for entity, relation, and triple extraction.
@@
Similarly, the experiments with different prompting styles showed that both simple inference prompts (predict) as well as advanced prompting with a Chain-of-Thought (Cot) rationale or a pipelined Extract-Critique-Refine (ECR) prompt showed improvements when automatic prompt optimizations were applied. These observations suggest that the automatic optimizers can adapt and propose effective refinements in different prompting scenarios related to triple extraction. Analysis at the relation-type level showed that optimized prompts tend to retain or improve performance on the majority of relation-types present in the dataset, without losing performance on those relation-types that were already handled correctly by the baseline. This indicates that APO not only boosts overall macro-level metrics but also does so broadly across different relation types, making the improvement generalizable and not narrowly focused on a few types.
@@
When the schema-complexity (i.e., the number of allowed relation types) and input context length (1x, 5x, and 10x of the original input text of test cases) were considered, it was seen that automatic prompt optimization was relatively more beneficial in more challenging scenarios. When the schema became more complex with a larger number of relations (from 100 up to 800 relations), extraction difficulty increased, and both baseline and optimized prompt performance dropped in relation and triple F1. Nevertheless, the decline was much more significant for baseline prompts, implying that the optimized prompts were more resilient to the schema-complexity. Similarly, when tested with increasingly long texts (by concatenating multiple test phrases together), both baseline and optimized prompt performance deteriorated, especially on relation and triple extraction (with relation extraction recall dropping due to challenges in identifying all relevant relations in a large blob of text). Optimized prompts can mitigate these context-length performance losses by still outperforming the baseline performance metrics for the single test case (1x), even when the input was 10x. The more descriptive optimized prompts with a set of optimal few-shot demonstrations seem to scale to larger input lengths more gracefully.
@@
When tested with different LLMs, it was observed that the triple extraction performance was mostly dominated by the capabilities of the task execution LLM rather than the LLM used to generate the prompts. The results imply that the choice of LLM at inference time seems more critical than the choice of LLM for prompt generation. In our experiments, prompts generated using a larger model (Llama 3.3-70B) were reasonably robust when applied to a different model (e.g. a 14B parameter model) on triple extraction. For example, the best prompt from Llama-70B tested on a smaller model achieved nearly the same triple F1 as that smaller model’s own optimized prompt. This suggests that there is some degree of model-agnostic prompt quality in optimised prompts (typically a good prompt containing generally useful instructions).
@@
When tested with multiple datasets, our results demonstrated some limitations in cross-dataset prompt transferability. When we used a different dataset’s training examples to optimize prompts (simulating a shift in domain or data distribution), the improvements became significantly less compared to using the examples from the same dataset. In particular, an optimized prompt trained on a secondary dataset yielded only marginal gains (on ${}1\%$ F1 increase) when tested on the target dataset’s text, compared to the much larger improvements (${}8\%$ or more) obtained when the training and testing data were from the same dataset. This raises a challenge for future work to improve the generalizability of automatic prompt optimizations.
The qualitative analysis of the generated prompts showed that the majority of the generated prompts were reasonably interpretable and aligned with human intuition in terms of explicitly providing details of the task, desired output format, and some guidance on how to perform the task.
@@
As future work, we are planning on exploring (a) cross-lingual prompt optimization where we plan on applying automatic prompt optimization techniques to multilingual settings, and (b) interactive human-in-the-loop prompt optimization, i.e. combining automatic prompt optimization with human feedback loops. An interactive automatic prompt optimization approach could allow a human analyst to guide the process by injecting domain knowledge or preferences (for example, selecting among automatically generated prompt candidates or refining them after a few iterations). Such approaches might fit well with agentic workflows with tool calling capabilities and decision making on when to get a human expert involved.
In conclusion, our empirical study demonstrates that automatic prompt optimization can be applied to triple extraction to improve knowledge graph construction, and also identifies some research directions for future explorations.

### v2

@@ 1 Introduction
Recent advances in large language models, which have significantly improved performance on core NLP tasks, have transformed the research area of KG construction from text [6, 7]. Most of the early large language model (LLM)-based methods approach the problem in two key ways. The first is as an extraction of factual and common-sense knowledge using pre-trained language models via prompt-based prediction of masked objects from partial sentences describing complete triples [8]. The second is by fine-tuning an LLM on the autoregressive generation of text for triple extraction [9, 10]. Similarly, in-context learning with decoder-only transformer models is applied for relation extraction [11]. Alternatives such as Re2G [12] and GraphRAG [13] use LLMs to build Retrieval Augmented Generation (RAG) systems [14].
@@ 2 Triple Extraction for Knowledge Graph Construction
To build a Knowledge Graph from text corpora, a system needs to convert the unstructured natural-language text into a set of elemental triples. Triple extraction involves sub-tasks such as identifying the core entities mentioned in the text (Named Entity Recognition or NER), along with identification/extraction of relations that constitute the facts mentioned in the text (Relation Extraction or RE). These sub-tasks can be pipelined, by performing entity extraction first, and then extracting relations and triples or in a joint manner, where entities, relations, and triples are extracted simultaneously (end-to-end relation extraction). In this paper, we focus on end-to-end triple extraction, where the input is a natural language text and the expected output is a set of triples.
@@ 3 Automated Prompt Optimization
Ma et al. [34] uses a strategy based on the greedy search. Gradient-based approaches such as ProTeGI [35], MAPO [36], TextGrad [37], etc. use gradient descent-like algorithms to optimize prompt embeddings according to a predefined performance objective [38]. Starting from a human prompt, these methods learn continuous valued embeddings or prompt proposals using the gradient directions. On the other hand, approaches such as InstructZero [39] optimize low-dimensional “soft prompts" using LLMs instead of directly optimizing discrete instructions. They balance exploitation vs exploration in this space using Bayesian Optimization techniques. In terms of discrete optimization, approaches such as EvoPrompt [40] circumvent access to gradients and internal parameters of LLMs using Evolutionary Algorithms to generate candidate prompts, thus making them more general. Another approach to the same problem leverages in-context learning. For example, Automated Prompt Engineer (APE) [41] generates prompts based on a few real-life examples, tailoring them to the task without needing extensive training data. APE treats the instruction as a program and optimizes it by searching through a pool of instruction candidates proposed by an LLM to maximize a chosen score function. A complementary approach to APO uses reinforcement learning-based strategies such as OIRL [42], which model the interaction between the query-prompt pair via a reward model for proposing and evaluating candidate prompts suited for arithmetic reasoning tasks. In contrast, meta-Prompting [43] uses structural and syntactical aspects of tasks to create general prompts that guide the generation of task-specific prompts. In lieu of template-based approaches, DSPy [44] frames prompt engineering as a declarative, compiler-driven optimization task. DSPy poses prompting as a computational graph pipeline for text transformation, where LLMs are invoked through parameterized self-improving modules. This provides great flexibility across diverse tasks given a computational budget.
@@ 4.1 Task
Let $\mathbf{x}=(x_{1},x_{2},\dots,x_{n})$ denote a passage of natural language text with a list of terms, and let $\mathcal{R}=\{r_{1},r_{2},\dots,r_{m}\}$ be a predefined set of canonical relation types. The goal of the triple extraction task is to identify: (a) A set of entity mentions $\mathcal{E}=\{e_{1},e_{2},\dots,e_{k}\}$ from the text, (b) a subset of relations $r\in\mathcal{R}$ expressed in the text, and (c) a set of directed triples $\bigl{\{}(e_{i},\,r,\,e_{j})\mid e_{i},e_{j}\in\mathcal{E},\;r\in\mathcal{R}\bigr{\}}$ that are represents the semantic relationships between the entity pairs that are mentioned in text. An example of the task, along with inputs and outputs, is illustrated in Figure 1.
@@ 4.3 Automatic Prompt Optimizers
DSPy [44] (Declarative Self-improving Python) is an open-source toolkit for modular prompt design as programming with structured and declarative natural-language modules. DSPy allows specifying input/output behaviour as a signature and selecting a module (e.g., predict, Chain-Of-Thought, ReAct) to assign a strategy for invoking the LLM. DSPy natives parse the LLM output based on the pre-defined signature. It allows for the optimization of prompts using a training set, which is analyzed to create a ‘sub-dataset’ and task descriptions, propose better prompts with LLM-generated instructions, and select effective few-shots examples. Specifically, we used MIPROv2 (Multiprompt Instruction PRoposal Optimizer Version 2) [47] to optimize instructions and few-shot examples simultaneously as a joint search problem. In the rest of the paper, when DPSy is mentioned, it is reffering to DSPy with MIPROv2. It first bootstraps a set of few-shot example candidates, proposes instructions grounded in different task dynamics, and finds an optimized combination of these joint candidates using Bayesian optimization. Figure 2 shows an example of an automatically optimized prompt using DSPy.
@@
APE [41] APE does not require any initial prompt. Given some examples of input and expected output, it generates several task instruction candidates, executes them using the target model, and selects the most appropriate instruction based on evaluation scores. By default, APE uses GPT-3 (text-davinci-002 to be more specific). We replace it with Granite-3.2-8B-Instruct, given its open-source availability. We used the forward generation template settings and randomly picked 10 examples from the SynthIE validation data to provide as few shots, and selected the generated prompt with the highest score based on the validation set of SynthIE.
@@ 5.1 Prompting strategy
Table 2 outlines the performance improvements in terms of macro Precision, Recall, and F1 obtained by prompt optimization. We observe that all prompting strategies result in a performance gain, as seen in Table 2. Of these, Chain-Of-Thought asked to solely extract triples obtains the best triple extraction F1 of $0.73$. The highest overall performance improvement is seen in the case of the Extract, Critique, and Refine pipeline run on entities, relations, and triples i.e., ECR (E-R-T) output, with a $16\%$ improvement in triple extraction F1.
@@ 5.3 Schema Complexity
To analyze this research question, we change the number of allowed relations that we include in the prompt. When a larger number of relations are included in the input text, we expect that the increased search space makes the task slightly harder for the systems with the increase in relation types. We examine eight settings with the number of allowed relation types ranging from 100800100-800.
@@ 5.5 Train, Validation and Test Datasets
As shown in Table 6, the results demonstrate that using the same dataset for both few-shot selection and prompt validation significantly outperforms the case when distinct datasets are used for training/validation and testing. For instance, when SynthIE is used as the test set, the Predict (E-R-T) prompt improves triple F1 by +0.01+0.01 when trained on the REBEL dataset, compared to a +0.08+0.08 improvement when the same dataset is used for both training and testing. Similarly, with Predict (T) and SynthIE as the test set, using REBEL for training/validation negatively impacts the optimized prompt, reducing triple F1 by 0.04-0.04 points. When REBEL is used as the test set, a similar trend is observed, with a larger margin of improvement when the same dataset is used for both training and testing compared to when different datasets are employed. Our hypothesis is that using few-shot examples and dataset summaries generated from a dataset that differs from the final test introduces some bias, leading to a prompt that is suboptimal for the test set.
@@ 6 Qualitative Analysis of Generated Prompts
Interesting, both the prompts yieleded similar results (when fed into the Llama3.3-70B) on the validation data - F1-Entities: 0.950.95/F1-Relations: 0.770.77/F1-Triples: $0.70$.
@@ 7 Conclusions
In this empirical study, we explored the application of automatic prompt optimization for triple extraction via extensive experimental benchmarking. We evaluated different settings by changing (a) the prompting strategy, (b) the LLM being used for prompt optimization and task execution, (c) number of canonical relations in the schema (schema complexity), (d) the length and diversity of input text, (e) the metric used to drive the prompt optimization, and (f) the dataset being used for training and testing. Based on the empirical results, automatic prompt optimization is a viable strategy for generating reasonable prompts for the triple extraction task, providing robust improvements against baseline prompts. The three prompt optimization strategies examine, namely, DSPy, APE, and TextGrad, consistently outperformed the baseline prompts in precision, recall, and F1 metrics for entity, relation, and triple extraction. Additionally, qualitative analysis of the generated prompts showed that a majority of the generated prompts were reasonably interpretable and in alignment with human intuition. Specifically, they provide explicit details of the task, desired output format, and some guidance on how to perform the task.
@@
The experiments with different prompting styles showed that both simple inference prompts (predict) as well as advanced prompting with a Chain-of-Thought (Cot) rationale or a pipelined Extract-Critique-Refine (ECR) prompts provided improvements during prompt optimizations. These observations suggest that the automatic optimizers can adapt and propose effective refinements across the board in scenarios related to triple extraction. Analysis at the relation-type level showed that optimized prompts tend to retain or improve performance on the majority of relation-types present in the dataset, without losing performance on those relation-types that were already handled correctly by the baseline. This indicates that APO not only boosts overall macro-level metrics but also does so broadly across different relation types, suggesting that the improvements are generalizable, i.e. do not narrowly focus on a few types.
@@
Experiments in increased schema-complexity (i.e., the number of allowed relation types) and input context length (1x, 5x, and 10x of the original input text) showed that automatic prompt optimization was considerably more beneficial the more challenging the scenario. When the schema became more complex with a larger number of relations (from 100 up to 800 relations), extraction difficulty increased, and both baseline and optimized prompt performance drops in terms of relation and triple F1. Nevertheless, the decline was relatively more significant for baseline prompts, implying that the optimized prompts were more resilient to increased schema-complexity. Increasingly long texts (by concatenating multiple test phrases together), both baseline and optimized prompt performance deteriorated, especially on relation and triple extraction (with relation extraction recall dropping due to challenges in identifying all relevant relations in a large blob of text). Optimized prompts can mitigate these context-length performance losses by still outperforming the baseline performance metrics for the single test case 1x all the way up to 10x. Descriptive optimized prompts with the right set of few-shot demonstrations seem to scale to larger input lengths quite gracefully.
Limitations and Challenges
@@
When tested with different LLMs, it was observed that the triple extraction performance was mostly dominated by the capabilities of the task execution LLM rather than the LLM used to generate the prompts. The results imply that the choice of LLM at inference time seems more critical than the choice of LLM for prompt generation. In our experiments, prompts generated using a larger model (Llama 3.3-70B) were reasonably robust when applied to a different model (e.g. a 14B parameter model) on triple extraction. For example, the best prompt from Llama-70B tested on a smaller model achieved nearly the same triple F1 as the smaller model’s own optimized prompt. This suggests that there is some degree of “model-agnostic prompt quality" maintained in the optimised prompts (typically a good prompt containing precise instructions that are generally useful).
@@
When tested with multiple datasets, our results also highlight some limitations in cross-dataset prompt transferability. For example, when we used a different dataset’s training examples to optimize prompts (this simulates a shift in domain/data distribution), the improvements reduced significantly compared to using the examples from the same dataset. In particular, an optimized prompt trained on a secondary dataset yielded only marginal gains (on $\sim 1\%$ F1 increase) when tested on the target dataset’s text, compared to the much larger improvements ($\sim 8\%$ or more) obtained when the training and testing data were from the same dataset. This highlights a set challenges for future work in prompt optimization.
Future Work
@@
We plan on exploring (a) cross-lingual prompt [48] optimization, where we will apply automatic prompt optimization techniques to multilingual settings, and (b) interactive human-in-the-loop prompt optimization, i.e. combining automatic prompt optimization with human feedback loops. An interactive APO approach could allow a human analyst to guide the process by injecting domain knowledge or preferences (for example, selecting among automatically generated prompt candidates or refining them after a few iterations). Such approaches might fit well with agentic workflows that integrate tool-calling capabilities and decision making on when to get a human expert involved.